OpenStudy (ja1):

Okay so I had this question and I am so confuso!!! Compare and contrast the two quadratic equations below. In order to receive full credit, use complete sentences to describe the following: The direction each parabola opens The vertex of each parabola y = x2 − 2x y = −2x2 + 4x − 3 I don't want you to help me with the question I need to know if the first equation is right?

OpenStudy (ja1):

Because I thought a quadratic equation had to have 3 terms not 2

OpenStudy (anonymous):

If we talk about a parabola generally, say in the form \(y = ax^2 + bx + c\), the sign of \(a\) (i.e., positive or negative) determines whether it is concave up or concave down. Also, the vertex of such a parabola is its intersection with the vertical line \(x=-\dfrac{b}{2a}\).

OpenStudy (ja1):

Yes but the first equation only has 2 terms not 3

OpenStudy (anonymous):

A quadratic equation can most certainly have two terms. It need only be a polynomial in \(x\) such that its highest power is \(x^2\) to be quadratic.

OpenStudy (anonymous):

@JA1 - that just means that \(b\) or \(c\) is zero.

OpenStudy (ja1):

So y = x2 − 2x is a Quadratic?

OpenStudy (anonymous):

Most certainly!

OpenStudy (abb0t):